A nut store usually sells cashews for $4.60 per pound and peanuts for $2.20 per pound. At the end of the month, the peanuts had not sold well. In order to sell 36 pounds of peanuts, the manager decided to mix the 60 lb of peanuts with some cashews and sell the mixture for $2.80 per pound. How many pounds of cashews should be mixed with the peanuts to ensure no change in the profit? (the value remains the same)
Answer by mananth(16946)

---------------- Price ---------------- quantity
peanuts 2.2 ---------------- 60 pounds
cashews 4.6 ---------------- x pounds
Mixture 2.8 ---------------- 60 + x pounds

2.2*60+4.6x=2.8 (60+x)
132+4.6x=168+2.8x
4.6x-2.8x=168-132
1.8 x = 36
/ 1.8
x = 20
20 pounds of 4.6 % cashews